OpenStudy (anonymous):

For example, if a jar contains 13 red marbles and 7 green marbles, the probability that a marble selected from the jar at random will be green is show some steps please

OpenStudy (rulnick):

Probability selected marble is green = # green / total # of marbles. So P(green) = 7 / (7+13) = 7/20.

OpenStudy (rulnick):

Or 0.35 if you prefer decimal.

OpenStudy (anonymous):

great ty

OpenStudy (anonymous):

total possible outcomes = 13 + 7 = 20 favourable for green = 7 so P(green) = 7/20

OpenStudy (rulnick):

You're welcome. Good luck. Bye.